# Data-Driven Machine Learning Approaches for Predicting In-Hospital Sepsis Mortality

### v1

@@ Abstract
Background
Sepsis is a severe condition responsible for many deaths in the United States and worldwide. Accurate prediction of sepsis outcomes is crucial for timely and effective treatment. Although previous studies have used machine learning to forecast outcomes, they faced limitations in feature selection and model comprehensibility, resulting in less effective predictions. Thus, this research aims to develop an interpretable and accurate Machine Learning model to help clinical professionals predict in-hospital sepsis mortality.
Methods
We analyzed ICU patient records from the MIMIC-III database based on specific criteria and extracted relevant data. Our feature selection process included a literature review, clinical input refinement, and using Random Forest to select the top 35 features. We performed data preprocessing, including cleaning, imputation, standardization, and applied the Synthetic Minority Over-sampling Technique (SMOTE) for oversampling to address imbalance, resulting in 4,683 patients, with admission counts of 17,429. We developed and compared the performance of Random Forest, Gradient Boosting, Logistic Regression, Support Vector Machine (SVM), and K-Nearest Neighbor (KNN) models.
Results
The Random Forest model was the most effective in predicting sepsis-related in-hospital mortality. It outperformed other models, achieving an accuracy of 0.90 and an AUROC of 0.97, significantly better than the existing literature. Our meticulous feature selection contributed to the model’s precision and identified critical determinants of sepsis mortality. These results underscore the pivotal role of data-driven machine learning in healthcare, especially for predicting in-hospital mortality due to sepsis, which greatly impacts patient outcomes and critical care.
Conclusion
This study represents a significant advancement in predicting in-hospital sepsis mortality, highlighting the immense potential of machine learning in healthcare. The implications are profound, offering a data-driven approach that enhances decision-making in patient care and reduces in-hospital sepsis mortality.
Introduction
@@
Fig 1: Patient Selection Process Graphical representation of patient inclusion criteria
@@
Fig 2: Data Extraction Process Graphical representation of health data and lab indicators Extraction
@@
Fig 3: Aggregation and Pivoting Process Graphical representation of data aggregation and pivoting for patient features
Data preprocessing
@@
Feature selection and feature importance
@@
Fig 4: Feature Importance Graphical Representation of the Top 35 Most Important Features Identified by a Random Forest
{adjustwidth}
0in-2.25in Age Lactate max Lactate avg Phosphate max Creatine Kinase min Albumin avg pH min Temperature avg Creatine Kinase median Phosphate avg Hematocrit max Albumin max Lactate median Temperature min Potassium max Bilirubin, Total avg Hematocrit min Creatine Kinase avg Albumin median Sodium avg Anion Gap max Neutrophils median Magnesium avg Hematocrit avg Neutrophils max Lactate Dehydrogenase max Temperature median Potassium, Whole Blood avg Lactate min Hematocrit median Temperature avg Creatine Kinase median Phosphate avg Hematocrit max Albumin max
@@
Model development and optimization
@@
Fig 5: Data Processing and Model Training Workflow Graphical representation of the steps from data extraction to the model prediction
Statistical analysis between cohors
@@ Results
{adjustwidth}
0in0.12in Characteristics Train Cohort Test Cohort T-Stat P-Value patient_age 64.152(15.196) 63.918(15.390) 0.594 0.552 Lactate_max 5.054(3.748) 5.010(3.767) 0.453 0.650 Phosphate_max 6.454(2.761) 6.416(2.676) 0.556 0.578 Lactate_avg 2.455(1.729) 2.425(1.754) 0.680 0.497 pH_min 7.205(0.184) 7.213(0.131) -2.024 0.043 Creatine_kinase_min 129.529(1583.876) 91.565(372.818) 1.729 0.084 Temperature_avg 37.042(0.654) 37.038(0.639) 0.247 0.805 Hematocrit_max 40.020(5.679) 40.045(5.671) -0.173 0.863 Anion_gap2_avg 14.205(2.979) 14.123(3.031) 1.055 0.292 Phosphate_avg 3.642(0.948) 3.610(0.962) 1.276 0.202 Albumin_avg 2.969(0.539) 2.952(0.535) 1.197 0.231 Potassium_max 5.981(1.404) 6.026(1.440) -1.227 0.220 Potassium_avg 4.143(0.358) 4.138(0.350) 0.583 0.560 Albumin_max 3.630(0.717) 3.630(0.731) -0.019 0.985 Temperature_min 36.020(1.152) 36.011(1.224) 0.301 0.763 Creatine_kinase_median 358.224(2934.465) 261.064(979.051) 2.234 0.026 Hematocrit_min 23.178(4.886) 23.017(4.770) 1.314 0.189 Bilirubin_total_avg 1.966(3.978) 1.794(3.596) 1.811 0.070 Temperature_median 37.034(0.672) 37.028(0.651) 0.342 0.732 Lactate_median 2.195(1.732) 2.174(1.774) 0.467 0.640 Albumin_median 2.944(0.576) 2.923(0.567) 1.406 0.160 Creatine_kinase_avg 450.048(3063.969) 331.290(1133.093) 2.543 0.011 Hematocrit_avg 30.597(3.477) 30.469(3.422) 1.460 0.144 Magnesium_avg 2.021(0.215) 2.013(0.207) 1.583 0.113 Sodium_avg 138.650(3.465) 138.588(3.271) 0.729 0.466 Phosphate_median 3.532(0.943) 3.502(0.963) 1.206 0.228 Lactate_min 1.156(0.997) 1.145(1.060) 0.391 0.696 Neutrophils_max 88.851(9.514) 89.436(8.341) -2.636 0.008 pH_avg 7.364(0.064) 7.366(0.062) -0.933 0.351 Anion_gap2_max 22.469(6.259) 22.444(6.382) 0.153 0.879 Sodium_max 146.455(5.418) 146.414(5.034) 0.306 0.760 Neutrophils_avg 75.895(12.170) 76.174(11.405) -0.937 0.349 Neutrophils_median 76.672(13.004) 76.833(12.391) -0.498 0.619 Hematocrit_median 30.282(3.564) 30.139(3.506) 1.583 0.114 Neutrophils_min 59.057(20.488) 58.880(20.251) 0.340 0.734
@@
Evaluation metrics proposed and baseline models’ performance
@@
{adjustwidth}
0in0.43in Model AUROC (95% CI) Accuracy Precision Recall F-Score RF 0.9715 [0.9656 - 0.9769] 0.9003 0.93 0.91 0.92 GB 0.8652 [0.8484 - 0.8808] 0.7901 0.87 0.79 0.83 LR 0.7701 [0.7491 - 0.7899] 0.7007 0.81 0.69 0.75 KNN 0.8840 [0.8687 - 0.8990] 0.7793 0.90 0.74 0.81 SVM 0.8628 [0.8453 - 0.8790] 0.7827 0.87 0.77 0.82
@@
Fig 6: Model Comparison The AUROC curves and scores for five models: Logistic Regression, Gradiant Boosting, Random Forest, SVM, and KNN.
@@
Fig 7: Model Comparison The bootstrap AUC scores for five models: Logistic Regression, Gradient Boosting, Random Forest, SVM, and KNN. The box plot shows the distribution of AUC scores obtained from bootstrap sampling.
Shapley Value analysis
@@
Fig 8: SHAP Analysis Summary plot for top feature impacts in sepsis prediction based on the Random Forest model.
Fig 9: SHAP Analysis Bar plot of mean absolute SHAP values, demonstrating the impact of each feature on the model’s prediction
Discussion
Summary of existing model compilation
@@ Discussion
Yong, L., Zhenzhou, L. This Study
@@
Study limitations and future research
@@ Conclusion
Acknowledgments
The authors would like to thank the developers of the MIMIC-III database for providing the extensive and detailed dataset on which our analysis was based.

### v2

@@ Abstract
Sepsis is a severe condition responsible for many deaths in the United States and worldwide, making accurate prediction of outcomes crucial for timely and effective treatment. Previous studies employing machine learning faced limitations in feature selection and model interpretability, reducing their clinical applicability. This research aimed to develop an interpretable and accurate machine learning model to predict in-hospital sepsis mortality, addressing these gaps. Using ICU patient records from the MIMIC-III database, we extracted relevant data through a combination of literature review, clinical input refinement, and Random Forest-based feature selection, identifying the top 35 features. Data preprocessing included cleaning, imputation, standardization, and applying the Synthetic Minority Over-sampling Technique (SMOTE) to address class imbalance, resulting in a dataset of 4,683 patients with 17,429 admissions. Five models—Random Forest, Gradient Boosting, Logistic Regression, Support Vector Machine, and K-Nearest Neighbor—were developed and evaluated. The Random Forest model demonstrated the best performance, achieving an accuracy of 0.90, AUROC of 0.97, precision of 0.93, recall of 0.91, and F1-score of 0.92. These findings underscore the potential of data-driven machine learning approaches to improve critical care, offering clinicians a powerful tool for predicting in-hospital sepsis mortality and enhancing patient outcomes.
keywords:
Sepsis, Critical Care, Mortality Prediction, MIMIC-III Database, Machine Learning
1 Introduction
@@ 2 Methodology
Figure 1: Patient Selection Process Graphical representation of patient inclusion criteria.
@@
Figure 2: Data Extraction Process Graphical representation of health data and lab indicators Extraction
@@
Figure 3: Aggregation and Pivoting Process Graphical representation of data aggregation and pivoting for patient features
2.1 Data preprocessing
@@
2.2 Feature selection and feature importance
@@
Figure 4: Feature Importance Graphical Representation of the Top 35 Most Important Features Identified by a Random Forest
@@
Age Lactate max Lactate avg
Phosphate max Creatine Kinase min Albumin avg
pH min Temperature avg Creatine Kinase median
Phosphate avg Hematocrit max Albumin max
Lactate median Temperature min Potassium max
Bilirubin, Total avg Hematocrit min Creatine Kinase avg
Albumin median Sodium avg Anion Gap max
Neutrophils median Magnesium avg Hematocrit avg
Neutrophils max Lactate Dehydrogenase max Temperature median
Potassium, Whole Blood avg Lactate min Hematocrit median
Temperature avg Creatine Kinase median Phosphate avg
Hematocrit max Albumin max
2.3 Model development and optimization
@@
Figure 5: Data Processing and Model Training Workflow Graphical representation of the steps from data extraction to the model prediction
2.4 Statistical analysis between cohors
@@ 3 Results
Characteristics Train Cohort Test Cohort T-Stat P-Value
patient_age 64.152(15.196) 63.918(15.390) 0.594 0.552
Lactate_max 5.054(3.748) 5.010(3.767) 0.453 0.650
Phosphate_max 6.454(2.761) 6.416(2.676) 0.556 0.578
Lactate_avg 2.455(1.729) 2.425(1.754) 0.680 0.497
pH_min 7.205(0.184) 7.213(0.131) -2.024 0.043
Creatine_kinase_min 129.529(1583.876) 91.565(372.818) 1.729 0.084
Temperature_avg 37.042(0.654) 37.038(0.639) 0.247 0.805
Hematocrit_max 40.020(5.679) 40.045(5.671) -0.173 0.863
Anion_gap2_avg 14.205(2.979) 14.123(3.031) 1.055 0.292
Phosphate_avg 3.642(0.948) 3.610(0.962) 1.276 0.202
Albumin_avg 2.969(0.539) 2.952(0.535) 1.197 0.231
Potassium_max 5.981(1.404) 6.026(1.440) -1.227 0.220
Potassium_avg 4.143(0.358) 4.138(0.350) 0.583 0.560
Albumin_max 3.630(0.717) 3.630(0.731) -0.019 0.985
Temperature_min 36.020(1.152) 36.011(1.224) 0.301 0.763
Creatine_kinase_median 358.224(2934.465) 261.064(979.051) 2.234 0.026
Hematocrit_min 23.178(4.886) 23.017(4.770) 1.314 0.189
Bilirubin_total_avg 1.966(3.978) 1.794(3.596) 1.811 0.070
Temperature_median 37.034(0.672) 37.028(0.651) 0.342 0.732
Lactate_median 2.195(1.732) 2.174(1.774) 0.467 0.640
Albumin_median 2.944(0.576) 2.923(0.567) 1.406 0.160
Creatine_kinase_avg 450.048(3063.969) 331.290(1133.093) 2.543 0.011
Hematocrit_avg 30.597(3.477) 30.469(3.422) 1.460 0.144
Magnesium_avg 2.021(0.215) 2.013(0.207) 1.583 0.113
Sodium_avg 138.650(3.465) 138.588(3.271) 0.729 0.466
Phosphate_median 3.532(0.943) 3.502(0.963) 1.206 0.228
Lactate_min 1.156(0.997) 1.145(1.060) 0.391 0.696
Neutrophils_max 88.851(9.514) 89.436(8.341) -2.636 0.008
pH_avg 7.364(0.064) 7.366(0.062) -0.933 0.351
Anion_gap2_max 22.469(6.259) 22.444(6.382) 0.153 0.879
Sodium_max 146.455(5.418) 146.414(5.034) 0.306 0.760
Neutrophils_avg 75.895(12.170) 76.174(11.405) -0.937 0.349
Neutrophils_median 76.672(13.004) 76.833(12.391) -0.498 0.619
Hematocrit_median 30.282(3.564) 30.139(3.506) 1.583 0.114
Neutrophils_min 59.057(20.488) 58.880(20.251) 0.340 0.734
3.1 Evaluation metrics proposed and baseline models’ performance
@@
Model AUROC (95% CI) Accuracy Precision Recall F-Score
RF 0.9715 [0.9656 - 0.9769] 0.9003 0.93 0.91 0.92
GB 0.8652 [0.8484 - 0.8808] 0.7901 0.87 0.79 0.83
LR 0.7701 [0.7491 - 0.7899] 0.7007 0.81 0.69 0.75
KNN 0.8840 [0.8687 - 0.8990] 0.7793 0.90 0.74 0.81
SVM 0.8628 [0.8453 - 0.8790] 0.7827 0.87 0.77 0.82
@@
Figure 6: Model Comparison The AUROC curves and scores for five models: Logistic Regression, Gradiant Boosting, Random Forest, SVM, and KNN.
@@
Figure 7: Model Comparison The bootstrap AUC scores for five models: Logistic Regression, Gradient Boosting, Random Forest, SVM, and KNN. The box plot shows the distribution of AUC scores obtained from bootstrap sampling.
3.2 Shapley Value analysis
@@
Figure 8: SHAP Analysis Summary plot for top feature impacts in sepsis prediction based on the Random Forest model.
Figure 9: SHAP Analysis Bar plot of mean absolute SHAP values, demonstrating the impact of each feature on the model’s prediction
4 Discussion
4.1 Summary of existing model compilation
@@
Yong, L., Zhenzhou, L. [17] This Study
@@
4.2 Study limitations and future research